# Approaching a dynamical extreme black hole horizon

### v1

@@ Abstract
We give an explicit closed form description of the late-time near-horizon approach to dynamical extreme Reissner–Nordström (DERN) black holes. These are spherically symmetric dynamical solutions of Einstein-Maxwell theory coupled to a neutral scalar that feature: (i) a spacetime metric which tends to that of a static extreme Reissner–Nordström (RN), and (ii) a scalar field which displays the linear Aretakis instability ad infinitum in the non-linear theory. We employ the two-dimensional Jackiw-Teitelboim (JT) gravity to solve explicitly for the non-linear s-wave dynamics of the four-dimensional theory near an ${\rm AdS}_{2}\times{\rm S}^{2}$ throat. For a teleologically defined black hole horizon, we impose boundary conditions on JT’s dilaton field (which encodes the gravitational dynamics) and the scalar matter as follows: (i) the JT dilaton decays at late times on the ${\rm AdS}_{2}$ boundary to a value that corresponds to a static extreme RN in the exterior, and (ii) the scalar obeys boundary conditions characteristic of linear Aretakis behavior on ${\rm AdS}_{2}$. We ensure our DERN solutions are singularity-free and we note that our approach to DERN is accompanied by a final burst of outgoing scalar matter flux leaking out of the ${\rm AdS}_{2}$ throat. The boundary conditions we impose on the JT dilaton place its late-time boundary profile on the threshold of black hole formation with sub-extreme and super-extreme RN on either side of our DERNs.
@@ I Introduction
The harbinger of weak instabilities associated with extreme black holes was the celebrated Aretakis instability [17, 18] for linear perturbations. Following that, the pioneering numerical study in [19] argued that there exist dynamical extreme Reissner–Nordström (DERN) black holes. They are the focus of the present work. DERNs are spherically symmetric solutions in the non-linear theory of Einstein-Maxwell coupled to a neutral scalar, whose metric tends to that of the static extreme Reissner–Nordström (ERN), while the scalar displays the Aretakis instability ad infinitum. [^2] Recently it was shown that for a planar extreme horizon, with infinite area, such asymptotic behavior is the generic non-linear fate of the linear Aretakis instability [20]. It was recently established rigorously that, within spherical symmetry, DERN is stable in the non-linear theory [21]. Moreover, it is argued numerically in [19] and conjectured mathematically in [21] that DERN lies on a threshold of black hole formation from critical collapse in this theory. This is a different threshold from the familiar Choptuik threshold [22] of naked singularities.
@@ II Aretakis instability and DERN
Our conventions in this paper are such that Reissner–Nordström (RN) takes the form: [^3] Our units are geometrized with $G=c=1$ and while most of the literature considers an electrically charged RN, it is just as good and slightly more convenient for us to consider the magnetic version above.
@@ II.1 Linear Aretakis instability on ERN
When Aretakis set out to explore stability of ERN under linear perturbations by a massless neutral scalar field $\phi$, he discovered that while $\phi$ itself does indeed decay at late times everywhere on and outside the future horizon $\mathcal{H}^{+}$, its transverse derivatives, $\partial_{\perp}^{n}\phi$, do not all decay at late times on the horizon [17, 18]. Specifically, for generic initial data, the $\ell$’th harmonic $\phi_{\ell}$ has its first $\ell$ transverse derivatives decay, $\partial_{\perp}^{k}\phi|_{\mathcal{H}^{+}}\to 0$ for $k\leq l$, but the next derivative approaches a constant $\partial_{\perp}^{l+1}\phi|_{\mathcal{H}^{+}}\to H_{\ell}$ and higher derivatives blow up beginning with $\partial_{\perp}^{\ell+2}\phi|_{\mathcal{H}^{+}}\sim H_{\ell}\,v$, where $v$ is the Killing time along $\mathcal{H}^{+}$. Here $H_{\ell}$, called an Aretakis constant, is non-vanishing for generic initial data with support on $\mathcal{H}^{+}$. For scattering data, supported away from the horizon, the instability is weaker in the sense that one needs an extra derivative to get to blow up, but otherwise still present [23, 24]. Precise late-time asymptotics on and outside $\mathcal{H}^{+}$, including the leading-order coefficients in terms of initial data, have been given in [23, 25].
A plethora of analytic and numerical work in the physics and mathematics literature has established the linear Aretakis instability as a robust dynamical feature associated with any perturbation of an extreme black hole horizon. This includes massive and charged scalars as well as (coupled) gravitational and electromagnetic perturbations, of extreme RN as well as Kerr black holes, and in four as well as higher dimensions [23, 26, 27, 28, 29, 30, 31, 32, 33]
@@ II.2 The nonlinear fate of the instability and DERN
gravitational backreaction has been shown numerically to regulate the instability by turning it into a transient one. Specifically, Ref. [19] showed that, generically, backreacting a spherically symmetric massless neutral scalar on extreme RN produces a near-extreme RN with the transverse derivatives of the scalar growing on its horizon for a timescale inversely proportional to its surface gravity, before eventually decaying. More precisely, [19] considered spherically symmetric characteristic initial data on a pair of bifurcate null hypersurfaces $\Sigma_{1}\cup\Sigma_{2}$ such that the spacetime $(M_{i},Q)$ and scalar $\phi\sim\mathcal{O}(\epsilon)$ data give rise to a future event horizon $\mathcal{H}^{+}$ which is the ERN one when $\epsilon=0$ and $M_{i}=Q$. Here, $Q$ is the Maxwell charge, which is conserved in the theory (6), and $M_{i}$ is the initial Bondi mass. Two kinds of data, whose linearization yields a perturbative test scalar evolving on a fixed ERN (as described in the previous subsection), were considered: $M_{i}=Q+\mathcal{O}(\epsilon^{2})$ and $M_{i}=Q+\mathcal{O}(\epsilon)$. [^4] The scalar’s stress tensor couples to the metric at quadratic order, so even a first order metric perturbation with $M_{i}=Q+\mathcal{O}(\epsilon)$ simply increases the RN mass but doesn’t couple to the scalar at leading order. For generic initial data of the first kind, $M_{i}=Q+\mathcal{O}(\epsilon^{2})$, compactly supported near $\mathcal{H}^{+}$, the numerical evolution settles down to a near-extreme RN with surface gravity $\kappa=\mathcal{O}(\epsilon)$. Prior to that, for a long time of order $\mathcal{O}(1/\epsilon)$, the transverse derivatives of the scalar behave as in the linear Aretakis story of the previous subsection: $\partial_{\perp}\phi|_{\mathcal{H}^{+}}$ remains approximately constant for $\mathcal{O}(1/\epsilon)$ time (before beginning a slow decay $\sim e^{-\kappa v}$), while $\partial_{\perp}^{2}\phi|_{\mathcal{H}^{+}}$ grows for $\mathcal{O}(1/\epsilon)$ time to a finite maximum value (before beginning its slow decay): $\lim_{\epsilon\to 0}\,\textrm{max}\left(\partial_{\perp}^{2}\phi|_{\mathcal{H}^{+}}\right)\neq 0$. This means that the Aretakis instability survives in the non-linear theory, albeit, generically only as a transient phenomenon. For data of the second kind, $M_{i}=Q+\mathcal{O}(\epsilon^{2})$, or for scattering data, supported away from $\mathcal{H}^{+}$, similar but weaker behavior was observed, namely the final RN black hole has $\kappa=\mathcal{O}(\epsilon^{1/2})$ and one needs an extra derivative to get to the finite growth on the horizon: $\lim_{\epsilon\to 0}\,\textrm{max}\left(\partial_{\perp}^{3}\phi|_{\mathcal{H}^{+}}\right)\neq 0$. An account of generic transient Aretakis instability using the JT theory was given in [34].
@@ III.1 ${\rm AdS}_{2}$ from three limits of RN
◦ Limit 2: sub-ERN $\rightarrow$ Rinder ${\rm AdS}_{2}\times{\rm S}^{2}$. Depicted in panels (b) and (e) in Fig. 2.
@@
$L\,T=\arctan\left(t+{1\over r}\right)-\arctan\left(t-{1\over r}\right)\,,\quad{R\over L}={1\over 2}r+{1\over 2}r\left(t^{2}-{1\over r^{2}}\right)$ (19)
@@ III.2 Linear Aretakis from ${\rm AdS}_{2}$
$\sigma|_{r=r_{0}}=\,$ $f(\pi/2)+g(\pi/2)-\frac{f^{\prime}(\pi/2)+g^{\prime}(\pi/2)}{v}+\frac{2f^{\prime}(\pi/2)}{r_{0}v^{2}}+\frac{f^{\prime\prime}(\pi/2)+f^{\prime\prime}(\pi/2)}{2v^{2}}$ (26)
@@
It is worth emphasizing that we are not thinking of the above as a derivation of the Aretakis instability for a scalar on ERN, but rather as a derivation of the appropriate boundary conditions to impose in a purely ${\rm AdS}_{2}$ analysis with respect to the preferred Poincare horizon. That said, we do expect it is possible to derive these boundary conditions from a full ERN analysis. Such a derivation would deal with the following two related aspects: first, the ERN ingoing Eddington-Finkelstein coordinate differs from the pure ${\rm AdS}_{2}$ one by a logarithmic correction that is diverging near the horizon $\hat{v}\approx v+2\ln r$; second, the relevant ${\rm AdS}_{2}$ for capturing near-horizon ERN physics is not the pure one considered here but instead one that is corrected by the anabasis perturbation discussed in the previous subsection, Eq. (10). Including the anabasis correction ensures, in the language of [36], a connected ${\rm AdS}_{2}$ which maintains the necessary (minimal) information from the asymptotically flat ERN physics. Using a connected ${\rm AdS}_{2}$, called a near-${\rm AdS}_{2}$ in JT theory, is essential for any consistent non-linear calculation which takes into account backreaction.

### v2

@@ Abstract
We give an explicit closed form description of the late-time near-horizon approach to dynamical extreme Reissner–Nordström (DERN) black holes. These are spherically symmetric dynamical solutions of Einstein-Maxwell theory coupled to a neutral scalar that feature: (i) a spacetime metric which tends to that of a static extreme Reissner–Nordström (RN), and (ii) a scalar field which exhibits the linear Aretakis instability ad infinitum in the non-linear theory. We employ the two-dimensional Jackiw-Teitelboim (JT) gravity to solve explicitly for the non-linear s-wave dynamics of the four-dimensional theory near an ${\rm AdS}_{2}\times{\rm S}^{2}$ throat. For a teleologically defined black hole horizon, we impose boundary conditions on JT’s dilaton field (which encodes the gravitational dynamics) and the scalar matter as follows: (i) the JT dilaton decays at late times on the ${\rm AdS}_{2}$ boundary to a value that corresponds to a static extreme RN in the exterior, and (ii) the scalar obeys boundary conditions characteristic of linear Aretakis behavior on ${\rm AdS}_{2}$. We ensure our DERN solutions are singularity-free and we note that our approach to DERN is accompanied by a final burst of outgoing scalar matter flux leaking out of the ${\rm AdS}_{2}$ throat. The boundary conditions we impose on the JT dilaton place its late-time boundary profile on the threshold of black hole formation with sub-extreme and super-extreme RN on either side of our DERNs.
@@ I Introduction
The harbinger of weak instabilities associated with extreme black holes was the celebrated Aretakis instability [17, 18] for linear perturbations. Following that, the pioneering numerical study in [19] argued that there exist dynamical extreme Reissner–Nordström (DERN) black holes. They are the focus of the present work. DERNs are spherically symmetric solutions in the non-linear theory of Einstein-Maxwell coupled to a neutral scalar, whose metric tends to that of the static extreme Reissner–Nordström (ERN), while the scalar exhibits the Aretakis instability ad infinitum. [^2] Recently it was shown that for a planar extreme horizon, with infinite area, such asymptotic behavior is the generic non-linear fate of the linear Aretakis instability [20]. It was recently established rigorously that, within spherical symmetry, DERN is stable in the non-linear theory [21]. Moreover, it is argued numerically in [19] and conjectured mathematically in [21] that DERN lies on a threshold of black hole formation from critical collapse in this theory. This is a different threshold from the familiar Choptuik threshold [22] of naked singularities.
@@ II Aretakis instability and DERN
Our conventions in this paper are such that Reissner–Nordström (RN) takes the form: [^3] Our units are geometrized with $G=c=1$ and while most of the literature considers an electrically charged RN, we find it slightly more convenient to work with the magnetic solution.
@@ II.1 Linear Aretakis instability on ERN
When Aretakis set out to explore the stability of ERN under linear perturbations by a massless neutral scalar field $\phi$, he discovered that while $\phi$ itself does indeed decay at late times everywhere on and outside the future horizon $\mathcal{H}^{+}$, its transverse derivatives, $\partial_{\perp}^{n}\phi$, do not all decay at late times on the horizon [17, 18]. Specifically, for generic initial data, the $\ell$’th harmonic $\phi_{\ell}$ has its first $\ell$ transverse derivatives decay, $\partial_{\perp}^{k}\phi|_{\mathcal{H}^{+}}\to 0$ for $k\leq l$, but the next derivative approaches a constant $\partial_{\perp}^{l+1}\phi|_{\mathcal{H}^{+}}\to H_{\ell}$ and higher derivatives blow up beginning with $\partial_{\perp}^{\ell+2}\phi|_{\mathcal{H}^{+}}\sim H_{\ell}\,v$, where $v$ is the Killing time along $\mathcal{H}^{+}$. Here $H_{\ell}$, called an Aretakis constant, is non-vanishing for generic initial data with support on $\mathcal{H}^{+}$. For scattering data, supported away from the horizon, the instability is weaker in the sense that one needs an extra derivative to get to blow up, but otherwise still present [23, 24]. Precise late-time asymptotics on and outside $\mathcal{H}^{+}$, including the leading-order coefficients in terms of initial data, have been given in [23, 25].
A plethora of analytic and numerical work in the physics and mathematics literature has established the linear Aretakis instability as a robust dynamical feature associated with any perturbation of an extreme black hole horizon. This includes massive and charged scalars as well as (coupled) gravitational and electromagnetic perturbations, of extreme RN as well as Kerr black holes, and in four as well as higher dimensions [23, 26, 27, 28, 29, 30, 31, 32, 33].
@@ II.2 The nonlinear fate of the instability and DERN
gravitational backreaction has been shown numerically to regulate the instability by turning it into a transient one. Specifically, Ref. [19] showed that, generically, backreacting a spherically symmetric massless neutral scalar on extreme RN produces a near-extreme RN with the transverse derivatives of the scalar growing on its horizon for a timescale inversely proportional to its surface gravity, before eventually decaying. More precisely, [19] considered spherically symmetric characteristic initial data on a pair of bifurcate null hypersurfaces $\Sigma_{1}\cup\Sigma_{2}$ such that the spacetime $(M_{i},Q)$ and scalar $\phi\sim\mathcal{O}(\epsilon)$ data give rise to a future event horizon $\mathcal{H}^{+}$ which is the ERN one when $\epsilon=0$ and $M_{i}=Q$. Here, $Q$ is the Maxwell charge, which is conserved in the theory (6), and $M_{i}$ is the initial Bondi mass. Two kinds of data, whose linearization yields a perturbative test scalar evolving on a fixed ERN (as described in the previous subsection), were considered: $M_{i}=Q+\mathcal{O}(\epsilon^{2})$ and $M_{i}=Q+\mathcal{O}(\epsilon)$. [^4] The scalar’s stress tensor couples to the metric at quadratic order, so a first order metric perturbation with $M_{i}=Q+\mathcal{O}(\epsilon)$ simply increases the RN mass but doesn’t couple to the scalar at leading order. For generic initial data of the first kind, $M_{i}=Q+\mathcal{O}(\epsilon^{2})$, compactly supported near $\mathcal{H}^{+}$, the numerical evolution settles down to a near-extreme RN with surface gravity $\kappa=\mathcal{O}(\epsilon)$. Prior to that, for a long time of order $\mathcal{O}(1/\epsilon)$, the transverse derivatives of the scalar behave as in the linear Aretakis story of the previous subsection: $\partial_{\perp}\phi|_{\mathcal{H}^{+}}$ remains approximately constant for $\mathcal{O}(1/\epsilon)$ time (before beginning a slow decay $\sim e^{-\kappa v}$), while $\partial_{\perp}^{2}\phi|_{\mathcal{H}^{+}}$ grows for $\mathcal{O}(1/\epsilon)$ time to a finite maximum value (before beginning its slow decay): $\lim_{\epsilon\to 0}\,\textrm{max}\left(\partial_{\perp}^{2}\phi|_{\mathcal{H}^{+}}\right)\neq 0$. This means that the Aretakis instability survives in the non-linear theory, albeit, generically only as a transient phenomenon. For data of the second kind, $M_{i}=Q+\mathcal{O}(\epsilon)$, or for scattering data, supported away from $\mathcal{H}^{+}$, similar but weaker behavior was observed, namely the final RN black hole has $\kappa=\mathcal{O}(\epsilon^{1/2})$ and one needs an extra derivative to get to the finite growth on the horizon: $\lim_{\epsilon\to 0}\,\textrm{max}\left(\partial_{\perp}^{3}\phi|_{\mathcal{H}^{+}}\right)\neq 0$. An account of generic transient Aretakis instability using the JT theory was given in [34].
@@ III.1 ${\rm AdS}_{2}$ from three limits of RN
◦ Limit 2: sub-ERN $\rightarrow$ Rindler ${\rm AdS}_{2}\times{\rm S}^{2}$. Depicted in panels (b) and (e) in Fig. 2.
@@
$L\,T=\arctan\left(t+{1\over r}\right)+\arctan\left(t-{1\over r}\right)\,,\quad{R\over L}={1\over 2}r+{1\over 2}r\left(t^{2}-{1\over r^{2}}\right)$ (19)
@@ III.2 Linear Aretakis from ${\rm AdS}_{2}$
$\sigma|_{r=r_{0}}=\,$ $f(\pi/2)+g(\pi/2)-\frac{f^{\prime}(\pi/2)+g^{\prime}(\pi/2)}{v}+\frac{2f^{\prime}(\pi/2)}{r_{0}v^{2}}+\frac{f^{\prime\prime}(\pi/2)+g^{\prime\prime}(\pi/2)}{2v^{2}}$ (26)
@@
It is worth emphasizing that we are not thinking of the above as a derivation of the Aretakis instability for a scalar on ERN, but rather as a derivation of the appropriate boundary conditions to impose in a purely ${\rm AdS}_{2}$ analysis with respect to the preferred Poincare horizon. That said, we do expect it is possible to derive these boundary conditions from a full ERN analysis. Such a derivation would deal with the following two related aspects: first, the ERN ingoing Eddington-Finkelstein coordinate differs from the exact ${\rm AdS}_{2}$ one by a logarithmic correction that is diverging near the horizon $\hat{v}\approx v+2\ln r$; second, the relevant ${\rm AdS}_{2}$ for capturing near-horizon ERN physics is not the rigid one considered here but instead one that is corrected by the anabasis perturbation discussed in the previous subsection, Eq. (10). Including the anabasis correction ensures, in the language of [36], a connected ${\rm AdS}_{2}$ which maintains the necessary (minimal) information from the asymptotically flat ERN physics. Using a connected ${\rm AdS}_{2}$, called a near-${\rm AdS}_{2}$ in JT theory, is essential for any consistent non-linear calculation which takes into account backreaction.